(a) Imagine that you have placed a limit order to 100 shares of Sallisaw Tool at a price of $38, although the stock is currently selling for $41. Discuss the consequences, if any, of each of the following situations.

(i) The stock price drops to $39 per share two months before cancellation of the limit order.

(ii) The stock price drops to $38 per share

(iii) The minimum stock price achieved before cancellation of the limit order was $38.50. When the limit order was cancelled, the stock was selling for $47.50 per share.

(b) You have been researching a stock that you like, which is currently trading at $50 per share. You would like to buy the stock if it were a little less expensive –say, $47 per share.

You believe that the stock price will go to $d70 by year and then level off or decline. You decide to place a limit order to buy 100 shares of the stock at $47 and a limit order to sell it at $70. It turns out that you were right about the direction of the stock price, and it goes straight to $75. What is your current position?
Answer #1

a)
i) If the stock price drops to 39 then order would not be executed because it is 1 $ more than the limit order.
ii) If the Stock drops to 38 the order would be executed and the amount spent = 38*100 shares = $ 3800
iii) If the stock is above limit order the order will not be executed and Stock has to be bought at current price of 41 and sold at 47.50 after stop order lapses.
Profit achieved = (47.50-41)*100 = $650

b)
The actual price achieved 75 after initial price of 50. The price never dropped to 47 hence no shares are bought. However since sell limit had been set at 70 the order 100 share will be sold at 70 So Amount received = 70*100 = 7000
Since actual price is 75 the loss is (75-70)*100 = 500


Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well